A quality-conscious disk manufacturer wishes to know the fraction of disks his company makes which are defective. Step 1 of 2 : Suppose a sample of 863 floppy disks is drawn. Of these disks, 717 were not defective. Using the data, estimate the proportion of disks which are defective. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Step 1 of 2 : from given data number of defectives in the sample =863-717=146

hence proportion of disks which are defective =146/863=0.169

( please revert if there are other parts of this)